### v2

# Repurposing Pre-trained Video Diffusion Models for Event-based Video Interpolation

Abstract
Video Frame Interpolation aims to recover realistic missing frames between observed frames, generating a high-frame-rate video from a low-frame-rate video. However, without additional guidance, the large motion between frames makes this problem ill-posed. Event-based Video Frame Interpolation (EVFI) addresses this challenge by using sparse, high-temporal-resolution event measurements as motion guidance. This guidance allows EVFI methods to significantly outperform frame-only methods. However, to date, EVFI methods have relied on a limited set of paired event-frame training data, severely limiting their performance and generalization capabilities. In this work, we overcome the limited data challenge by adapting pre-trained video diffusion models trained on internet-scale datasets to EVFI. We experimentally validate our approach on real-world EVFI datasets, including a new one that we introduce. Our method outperforms existing methods and generalizes across cameras far better than existing approaches.
Figure 1: We compare our proposed approach RE-VDM, which adapts a pre-trained video diffusion model for event-based video frame interpolation on unseen real-world data, with three baselines: frame-only interpolation method GIMM-VFI-R-P [14], test-time optimization of a pre-trained video diffusion model via Time Reversal [11], and event-based interpolation method CBMNet-Large [21], trained on the same dataset as our method. The left-most column shows the start frame, end frame, and reference interpolation frames overlaid with events. Leveraging data priors in the pre-trained video diffusion model and the diffusion process, along with motion guidance controlled by input events, our approach demonstrates superior generalization performance on unseen real-world frames with substantial motion.
1 Introduction
Event cameras are a novel class of neuromorphic sensors offering unique advantages, including high dynamic range and high temporal resolution [12]. One significant application of event cameras is Event-based Video Frame Interpolation (EVFI) [41, 42, 51, 40, 28, 21]. By capturing traditional frames alongside high temporal resolution events, event data can fill in missing motion information between frames, aiding in interpolation. This approach avoids the ill-posed problem of frame-only interpolation, where large motion between frames introduces excessive degrees of freedom, making the interpolation of frames infeasible.
Current state-of-the-art (SOTA) EVFI methods achieve impressive performance when trained and tested on the same dataset. However, when the test set differs from the training set with changes in data and motion distribution, these methods experience significant performance degradation, as we can see in Fig. 1. This issue arises from two main factors. First, compared to broader fields like video generation, the EVFI field is much smaller, with limited data quantity and quality. Second, the models used in EVFI are highly specialized, with constrained representational power.
Recent advances in generative AI, particularly in video generation, have spurred significant progress in the field. Leading companies have invested substantial resources into building massive, high-quality datasets [8] and developing foundation models for video generation [17, 8]. Most of these foundation models are video diffusion models. By leveraging millions of high-quality commercial video clips and billions of parameters, these diffusion models achieve impressive generalization performance compared to previous approaches. A comparison between video generation and EVFI fields is presented in Table 1. The primary differences between these fields lie in dataset quality/size and model parameter scale, the EVFI field lacks the extensive commercial datasets and large model sizes found in video generation, impacting its models’ generalization and representational power.
This raises a natural question: Can we adapt pre-trained video diffusion foundation models for EVFI to leverage their learned data priors and model design advantages?
In this work, we introduce the first approach, RE-VDM, for adapting pre-trained video diffusion foundation models to the EVFI task, aiming to bridge EVFI with generative AI by leveraging robust data priors and foundation models. To achieve this, we address key technical challenges. First, for data-efficient adaptation, we train an additional subset of the model with event-based control on small EVFI datasets inspired by [49], while keeping the original weights frozen to avoid catastrophic forgetting of learned data priors as explained in Section 3.2. Second, while video generation aims to produce diverse videos without strict realism constraints, EVFI requires ground-truth fidelity to accurately recover missing frames. To preserve motion and appearance fidelity, we employ upsampling and Per-tile Denoising and Fusion as explained in Section 3.4 as the test-time optimization for video generation with high-fidelity appearance and event-based motion control. Finally, unlike video generation, which is an extrapolation task based on an image or text prompt, EVFI requires interpolation between the start and end frames. To address this, we develop a test-time optimization approach that generates event-based video latents from both the start and end frames at each denoising step, fusing them for a consistent, interpolated result that incorporates information from both frames, as explained in Section 3.5
For a robust evaluation, we compare the generalization performance of our approach against representative frame-only, event-based, and pre-trained video diffusion models for frame interpolation across multiple datasets, including self-collected Clear-Motion test sequences. The results demonstrate the potential and advantages of this adaptation. We also discuss the limitations observed, highlighting the exploratory nature of our work. To summarize, our contributions are
  • We present the first approach for adapting pre-trained video diffusion foundation models to the Event-based Video Frame Interpolation task, leveraging the strong data priors learned from large video generation datasets and the inherent advantages of video diffusion models.
  • Experimental results demonstrate the strong potential of the adapted video diffusion model for the EVFI task, showing excellent generalization on unseen real-world data and consistency in reconstructed frames.
Feature Video Generation EVFI
Dataset Size [7, 32, 23] $10^{7}$ $10^{2}$
Dataset Quality [7, 32, 23, 21, 42, 39] Commercial Custom
Model Type [7, 42, 40, 28, 21] Video Diffusion (Foundation) Models Custom Models
Model Size [7, 32, 42, 40, 28, 21] $10^{10}$ $10^{8}$
Video Generation Type [7, 42, 40, 28, 21] Extrapolation Interpolation
Table 1: Comparison between video generation and event-based video frame interpolation (EVFI) fields reveals major differences in dataset size, quality, and model design. Video generation datasets are orders of magnitude larger than EVFI datasets and are of commercial-level quality. Models for video generation are typically foundation models with billions of parameters. With abundant training data and large-scale foundation models, video diffusion models are well-suited for generalized video generation tasks.
2 Related Work
2.1 Event Camera and Video Frame Interpolation
Event cameras are neuromorphic sensors that capture the brightness changes in the scene, and the brightness changes are primarily due to the camera’s motion and object motion in the scene. For this reason, events captured by event cameras can be used to compute motion-related information, for example, optical flow, motion segmentation, and ego-motion [6, 4, 3, 5, 29, 15, 37, 55, 54, 47]. Various methods have been proposed to reconstruct intensity videos from event streams alone, demonstrating that while feasible, the quality of such reconstructions is significantly limited by camera and object motion [9, 33]. Combining events with images extends beyond event-only applications, offering powerful capabilities for tasks like image deblurring and video frame interpolation/extrapolation [41, 42, 51, 10, 40]. This approach leverages events as cues to bridge the gap between frames, thus avoiding the ill-posed challenges faced by frame-only methods [18, 31, 25, 20, 1, 22, 45, 53, 48, 14] when interpolating between frames with substantial motion.
2.2 Diffusion and Video Diffusion Models
Diffusion models [16, 38] is an emerging class of image generative models that model the reverse diffusion process of recovering data from noise. Starting from data $\mathbf{x}_{0}$ from the data distribution $p_{data}(\mathbf{x})$, The diffusion model tries to reverse the forward diffusion process as an $N$ step Markov chain $\{\mathbf{x_{i}}\},i\in[0,N]$ that satisfies the marginal distribution
$p_{\alpha_{i}}(\mathbf{x}_{i}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{i};{\alpha(t)}\mathbf{x}_{0},\sigma(t)^{2}\mathbf{I}).$ (1)
Here, $\alpha(t),\sigma(t)$ represents a noise schedule. Diffusion models have demonstrated remarkable image generation quality and diversity. Stable Diffusion [34] shifts the generation process to a low-dimensional latent space, greatly reducing computational cost. Trained on massive online text-image pair datasets [36], Stable Diffusion exhibits strong image generation capabilities. Video diffusion models [17, 8] add temporal layers on top of existing image diffusion models to jointly denoise multiple consistent image frames. Stable Video Diffusion (SVD) [8] is built on top of latent diffusion model and trained on massive video datasets, serving as a strong foundational model for video generation tasks.
2.3 Controllable Diffusion Generation Through Fine-Tuning
There have been numerous works on generation conditioned on additional user inputs or control [34, 35, 49, 46]. Stable Diffusion [34] uses cross-attention layers in the diffusion U-Net to inject text control. Dreambooth [35] showed that a diffusion model can be fine-tuned for conditional control. ControlNet [49] modifies a trainable copy of the original diffusion U-Net for conditional control. In addition to fine-tuning diffusion models for added controllability, various training-free methods have been proposed to alter the sampling process of image and video diffusion models for more flexible and generalized content synthesis [2, 11, 27, 52]. Latent diffusion models have limitations on generating a limited set of resolutions in a downsampled latent space. MultiDiffusion [2] provides a solution for resolution-agnostic generation by tile-based denoising. When combined with conditional control, MultiDiffusion [2] has shown to be effective in generating synthetic detailed images under the condition of fine-grained condition control [19, 43]. On the video generation side,  [11] proposes a test-time sampling method that empowers video diffusion models to control camera and object motion.  [52] enables consistent long-frame generation outside of the trained frame length of a video diffusion model.
3 Proposed Approach
3.1 Pipeline Overview
Figure 2 presents an overview of our pipeline RE-VDM, for adapting pre-trained video diffusion models to event-based video generation and interpolation tasks. The key difference is that interpolation applies our EVDS on both the start frame $I_{s}$ for forward-time video denoising/generation and the end frame $I_{e}$ for backward-time video denosing/generation. At each denoising step, forward-backward consistency is enforced to achieve accurate interpolation.
In this work, without loss of generality, we assume that the pre-trained video diffusion model used is a Latent Diffusion Model, as is common in most video diffusion models.
Figure 2: The overview of our proposed approach RE-VDM: for adapting pre-trained video diffusion models includes two tasks: event-based video generation and interpolation. For video generation, the method utilizes the start frame $I_{s}$ and forward-time events $E_{s}$. For interpolation, it incorporates both the start frame $I_{s}$ and forward-time events $E_{s}$, as well as the end frame $I_{e}$ and backward-time events $E_{e}$ to achieve consistent results. Unlike video generation, for interpolation, a denoising step $t$ concludes with Two-side Fusion instead of EVDS.
3.2 Data-Efficient Adaptation of Pre-trained Video Diffusion Models
In this section we will explain the details for introducing the additional event-based motion control to pre-trained video diffusion models. As discussed in Section 1 and shown in Table 1, the large training datasets and model size of pre-trained video diffusion foundation models provide impressive generalization power for video generation tasks. However, these factors also limit the fine-tuning flexibility of the pre-trained model weights on existing EVFI datasets due to the risk of catastrophic forgetting, stemming from significant differences in data size and quality.
To prevent catastrophic forgetting of the learned data priors in pre-trained video diffusion foundation models, we drew inspiration from the [49] approach for controllable image diffusion models, known for its impressive performance with small data adaptation. In our design, we keep the original weights of the video diffusion models frozen during training. Event-based control is introduced by training only a subset of blocks, copied from the original denoiser network, to serve as additional residuals to corresponding block outputs in the frozen denoiser network. A detailed illustration is provided in Figure 3.
Figure 3: The illustration depicts our training scheme for adapting a pre-trained video diffusion model to event-based video denoising. Our approach uses a frozen denoiser network from the pre-trained model, augmented with a trainable subset of blocks copied from the frozen denoiser.
The added event encoder and trainable copy aim to learn the correct encoding for events as the control condition, $E_{cond}$, as well as the block output residuals for the frozen pre-trained denoiser network by optimizing the trainable copy. We use the standard diffusion objective, which minimizes the mean squared error loss between the predicted noise $\epsilon^{\prime}$ and the ground truth noise $\epsilon$ added to the video latent for a uniformly randomly sampled denoising time $t$.
During inference, the only difference from the training scheme is the absence of a given video latent $Z_{t}$ and noise $\epsilon$ for a sampled denoising time $t$. Instead, the initial noisy latent $Z_{T}$ is generated from standard Gaussian noise $N(0,1)$ at the beginning of the denoising process.
3.3 Events as the Control Condition
To introduce events as the control condition for the video generation process, we incorporate two main components: 1) Event Representation and 2) Event Encoder. For event representation, we adopt a multi-stack approach, inspired by [30, 26], which captures both fast- and slow-moving objects within a multi-channel, frame-like format. An example of this multi-stack event representation is illustrated in Figure 4. Based on this representation, our event encoder consists of a series of convolutional layers with stride to downsample the input event representation into the event latent $E_{cond}$, matching the dimension of the image latent $I_{cond}$.
Figure 4: Our multi-stack event representation is illustrated as follows. The stack begins from the target frame at $t_{i}$ and expands backward in time to the previous frame $t_{i-1}$. Within each stack, the number of events accumulated from $t_{i}$ is halved from the previous stack. In the long stack ($m=0$), the slowest-moving objects, such as around the human head, appear sharp; in the middle stack ($m=1$), slower-moving objects, like the human arm, are clear; and in the short stack ($m=2$), the fastest-moving objects, such as the ball, are sharp. This approach ensures that the event data provides adequate control information for generating frame $t_{i}$.
Figure 5: The VAE encoding/decoding loss of small details in the original input image. In (a), we show the original image, sized 970 x 625. In (b), we pad the image to the nearest multiple of 8, then encode and decode it back. After decoding, the PSNR drops to 21.50, with noticeable detail loss in the zoomed-in view. In (c), we pad to the nearest multiple of 8, upsample to twice the original width and height, then encode and decode. After decoding, the PSNR increases to around 24.92, with no significant loss of details in the zoomed-in view.
3.4 Per-tile Denoising and Fusion
Due to the substantial memory demands of training and generating high-resolution video, most existing video diffusion models, such as Stable Video Diffusion (SVD) [7], are Latent Diffusion Models, meaning they operate in the latent space with downsampled spatial resolution rather than in the original pixel space. As illustrated in Figure 3, for an input RGB image/video of shape [$B$, $F$, 3, $H$, $W$], where $B$ is the batch size, $F$ is the number of frames, $W$ is the width, and $H$ is the height, a latent encoder ($\mathcal{E}$) and decoder ($\mathcal{D}$), typically based on a Variational Autoencoder (VAE), convert the input into image latents of shape [$B$, $F$, $C_{latent}$, $\frac{H}{d}$, $\frac{W}{d}$] and decode it back.
The Latent VAE encoder/decoder approach is effective for video generation, where realism is less critical, as the task often involves stylized animation rather than high-detail photorealism. However, for EVFI, minimizing the loss from downsampled encoding/decoding is essential, as our goal is to recover realistic missing frames between keyframes. A key observation is that small details can be lost after VAE encoding/decoding. Upsampling the input mitigates this loss, as shown in Figure 5, where increasing input upsampling effectively reduces encoding/decoding loss on fine details.
Additionally, as illustrated in Figure 3, spatial downsampling of the input image to create image latents can also reduce event control accuracy, since the trainable copy’s input consists of combined image latents $I_{cond}$ and event latents $E_{cond}$ of the same dimensions. However, upsampling input videos substantially increases computational costs for both training and testing. To address this, we divide the upsampled input video into overlapping cropped tiles of fixed width and height, applying EVDS to each tile independently. We then fuse the denoised latents of tiles to achieve consistent latents for each image and video, preserving shared context across overlapping regions
Our Per-tile Denoising and Fusion process is illustrated in Figure 6. We employ tiled diffusion [2] by first dividing the latent canvas into $n$ overlapping tiles. For the denoised tile latents at tile $i$ at denoising step $t$, we use an accumulation process to fuse them into the denoised latent for the entire upsampled video. The accumulation process takes in $n$ tile latents and outputs the fused latent:
$\tilde{Z}_{t-1}=\sum_{i=1}^{n}\frac{W_{i}}{\sum_{j=1}^{n}W_{j}}\otimes Z^{i}_{t-1},$ (2)
where $\tilde{Z}_{t-1}\in\mathbb{R}^{F\times H\times W\times d}$, with $F$ representing the number of frames, $H$ and $W$ the height and width of the upsampled latent canvas, and $d$ the latent dimension. $Z^{i}_{t-1}$ contains the denoised latents of the $i$-th grid tile, $Z^{i}_{t-1}\in\mathbb{R}^{F\times h\times w\times d}$, where $h$ and $w$ denote the height and width of each latent tile. $W_{i}\in\mathbb{R}^{F\times h\times w}$ represents pixel-wise weights for the $i$-th tile. In our implementation, we assign equal weights, averaging the denoised latents across overlapping tiles.
Figure 6: The Per-tile Denoising and Fusion process is a test-time optimization applied during inference to enhance the fidelity of video generation appearance and improve event-based motion control accuracy. During each denoising step, our model operates on upsampled tiles of the input image and event representations to predict denoised latents for each tile $i$ at a denoising time $t$ ($Z_{t-1}^{i}$). These denoised tile latents are then accumulated to obtain the predicted denoised latents for the entire video (${\tilde{Z}}_{t-1}$) at that denoising step $t$.
Method Category BS-ERGB (3 skips) HQF (3 skips) Clear-Motion (15 skips)
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Time Reversal [11] (ECCV’24) Video Diffusion 17.86 0.59 0.27 17.92 0.53 0.18 13.71 0.70 0.44
DynamiCrafter [46] (ECCV’24) Video Diffusion 15.47 0.54 0.40 15.99 0.56 0.35 11.42 0.62 0.52
RIFE [18] (ECCV’22) Frame-only VFI 23.30 0.83 0.09 25.34 0.76 0.05 16.93 0.76 0.35
PerVFI [45] (CVPR’24) Frame-only VFI 23.53 0.79 0.11 27.13 0.81 0.04 16.74 0.75 0.30
InterpAny-Clearer [53] (ECCV’24) Frame-only VFI 25.51 0.85 0.14 27.93 0.83 0.06 17.21 0.77 0.32
EMA-VFI [48] (CVPR’23) Frame-only VFI 26.01 0.86 0.18 28.31 0.84 0.10 17.79 0.78 0.44
GIMM-VFI-R-P [14] (NeurIPS’24) Frame-only VFI 25.56 0.84 0.08 27.96 0.83 0.04 17.77 0.77 0.26
CBMNet-Large [21] (CVPR’23) EVFI 26.24 0.79 0.22 28.73 0.84 0.09 22.26 0.86 0.39
RE-VDM (Ours) EVFI + Video Diffusion 27.74 0.88 0.12 29.04 0.89 0.06 22.94 0.88 0.37
Table 2: We compare our method on real-world EVFI datasets against representative and SOTA methods, including the pre-trained video diffusion model (Video Diffusion), frame-only methods (Frame-only VFI), and event-based video frame interpolation method (EVFI), where red indicates the best metric and blue indicates the second-best. For CBMNet-Large, we use publicly available model checkpoints trained on the same BS-ERGB dataset as our approach. For the BS-ERGB and HQF datasets, we use a skip number of 3 as in the EVFI papers, we evaluate all interpolated frames. For our Clear-Motion test sequences, we use a skip number of 15 for evaluation on large motion scenarios, we evaluate the 4th, 8th, and 12th interpolated frames. To control for VAE encoding/decoding loss, as discussed in Section 4.2, we apply VAE encoding/decoding to all model outputs.
3.5 From Video Generation to Frame Interpolation
Our approach thus far is designed to generate realistic video controlled by events from the start frame, suited for event-based video generation (extrapolation). However, the EVFI task requires frame interpolation, which involves utilizing information from both the start and end frames.
The most straightforward approach to achieve frame interpolation would be to fine-tune base video diffusion models to accept an additional channel for the second image. However, as previously discussed, fine-tuning pre-trained video diffusion foundation models on the EVFI dataset risks catastrophic forgetting of the learned data priors, undermining the purpose of adaptation.
To address this, we opted for a test-time optimization approach that avoids further training, enabling our video generation model to perform EVFI. As shown in Figure 2, inspired by [11], our key idea is to run the EVDS at each denoising step from both the start frame $I_{s}$ and the end frame $I_{e}$, producing two sets of predicted denoised video latents. We then apply a Two-side Fusion of these denoised latents to obtain consistent results. This approach allows us to achieve frame interpolation, effectively leveraging information from both frames. The details of the Two-Side Fusion algorithm are as follows: For a denoising time $t$, the forward denoised video latents from the start frame $I_{s}$ and forward-time events $E_{s}$ are $Z^{s}_{t-1}=EVDS(I_{s},E_{s})$, and the backward denoised video latents from the end frame $I_{e}$ and backward-time events are $Z^{e}_{t-1}=EVDS(I_{e},E_{e})$, with $Z^{s}_{t-1},Z^{e}_{t-1}\in\mathbb{R}^{F\times H\times W\times d}$. To make the final denoised video latent $\tilde{Z}_{t-1}$ consistent with information from both the start and end frames, we assign linearly increasing weights, $W_{f}\in\mathbb{R}^{F}$, where $W_{0}=0$ and $W_{1}=1$. This allows us to fuse the final video latents $\tilde{Z}_{t-1}$ as follows:
$\tilde{Z}_{t-1}=W_{f}\otimes Z^{s}_{t-1}+(1-W_{f})\otimes flip(Z^{e}_{t-1})$ (3)
where $flip$ denotes reversing the temporal order of the frames in $Z^{e}_{t-1}$. The intuition behind this approach is that frames near the start frame rely more heavily on information from the start frame and forward-time events, while frames near the end rely more on the end frame and backward-time events, in the middle, they balance information from both.
4 Experimental Results
4.1 Datasets and Implementation Details
To robustly evaluate the generalized frame interpolation performance of our approach, we select several real-world event camera datasets. The BS-ERGB dataset [41], the most cited for EVFI, provides aligned RGB and event video with large motion between frames, mainly featuring static cameras capturing dynamic objects. The HQF dataset [39] includes grayscale video aligned with events, with more sequences involving a moving event camera. In addition to existing datasets, we include our self-collected Clear-Motion test sequences, capturing both object and camera motion. Detailed information is provided in the Supplementary Material. Clear-Motion is a self-collected dataset used solely for testing, eliminating the possibility of data leakage. This makes it ideal for robustly evaluating the zero-shot generalization performance of all models. The pre-trained video diffusion model we used is Stable Video Diffusion [7] for 14-frame image-to-video generation. We provide detailed training information in the supplementary material.
4.2 Evaluation Strategy
To evaluate generalized frame interpolation performance, we compare our approach with representative methods from three categories: frame-only interpolation, pre-trained video diffusion models, and event-based interpolation. We use publicly available CBMNet checkpoints trained on the BS-ERGB dataset for evaluation. For Per-tile Denoising and Fusion in our method, which uses upsampled tiles to reduce detail loss and enhance event control accuracy, we apply an upsample factor of 2 for the BS-ERGB and Clear-Motion datasets and 3 for the HQF dataset.
For evaluation metrics, we use average frame PSNR [13], SSIM [44], and LPIPS [50] for the interpolated frames. Since SVD is a latent diffusion model that uses a VAE to encode and decode between latent and pixel space, we apply VAE encoding and decoding to all model outputs. This eliminates the impact of discrepancies between the VAE-decoded frame distribution and the ground truth frame distribution, accounting for differences in small details and tone mapping.
4.3 Evaluation on Video Frame Interpolation
In this section, we present both quantitative and qualitative results demonstrating the effectiveness of our method compared to representative video frame interpolation approaches, as shown in Table 2. Our method outperforms all recent SOTA VFI baselines in terms of PSNR and SSIM. Qualitative comparisons in Figures 7 and 8 further show that recent frame-only VFI methods struggle with large and complex motion due to the ill-posed nature of the task. This limitation also affects pre-trained video diffusion models, which often exhibit a bias toward animation-style generation and suffer from spatial-temporal detail loss.
In contrast, the SOTA event-based interpolation method, CBMNet-Large, can handle large motions by leveraging cues from events, similar to our approach. However, its performance on both the test set from the training dataset and unseen datasets is consistently lower than ours across all metrics. This difference is due to two factors: 1) as discussed in Section 1 and Table 1, the pre-trained video diffusion foundation model in our approach offers a superior data prior from its extensive training on large video generation datasets, and 2) our model’s backbone, a highly representative pre-trained video diffusion model, provides advantages over custom models typically used in EVFI.
The effectiveness of our approach on real-world, unseen data for video frame interpolation is clearly demonstrated in the qualitative comparisons in Figure 7. Due to the ill-posed nature of interpolating frames with large motion, frame-only methods, despite often achieving better LPIPS scores, struggle to accurately reconstruct object locations and appearance. While CBMNet-Large uses event data to guide interpolation and achieves similar object and texture placements, it lacks generalization on unseen object and camera motion compared to our method, as shown in Figure 7, despite being trained on the same dataset. Additionally, CBMNet-Large suffers from frame-to-frame inconsistencies due to its frame-by-frame inference design, evident in Figure 8. CBMNet-Large produces frames with inconsistent textures and shapes on the ball and hand, whereas our method accurately reconstructs the challenging, regular shapes of both. This advantage is due to the strong data priors and foundational design of the pre-trained video diffusion model, allowing for consistent reconstruction across frames. These results support our unique contributions and motivation in adapting pre-trained video diffusion models for EVFI.
4.4 Ablation Study
In the Supplementary Material, we provide quantitative and qualitative ablation studies on two key components for adapting pre-trained VDMs to EVFI: 1) Comparing video generation and video interpolation using Two-Side Fusion. 2) Ablating the effect of upsampling in Per-tile Denoising and Fusion. Both quantitative and qualitative results demonstrate the effectiveness of the two main components in our approach for adapting pre-trained VDMs to the EVFI task.
Figure 7: The qualitative comparison on the Clear-Motion sequence Texture_Box, featuring translation and rotation of a textured box, demonstrates that our method achieves the most consistent and accurate frame reconstruction. Other methods experience significant detail loss, artifacts, and difficulty maintaining regular object shapes, particularly in the characters and the bottom-right coconut.
Figure 8: The qualitative comparison on the BS-ERGB test sequences, featuring fast motion of a hand and a rising, spinning ball, shows that our method is the only approach capable of maintaining the regular texture shapes on both the ball and hand in this challenging scenario.
5 Limitations
Despite the promising results, adapting pre-trained video diffusion models to the EVFI task presents several limitations. First, the pre-trained VAE encoder and decoder, used for conversion between pixel space and latent space, have limited representational power, often resulting in color shifts and loss of small details. Second, the adapted video diffusion model tends to smooth out certain regions, likely due to its difficulty in preserving fine textures within the downsampled latent space or the need for a prohibitively high upsampling factor to prevent smoothing. Finally, the diffusion-based methods are more memory-intensive and time-consuming than other methods, we provide the detailed model run time, memory comparison in the Supplementary Material. These limitations stem from the current state of pre-trained video diffusion foundation models. However, as the field of video generation evolves rapidly, we anticipate that many of these issues could be mitigated or resolved with newer, more advanced pre-trained models in the near future.
6 Conclusion
In this work, we explore, for the first time, the adaptation of pre-trained video diffusion foundation models for the event-based video frame interpolation (EVFI) task, as a foundational step toward integrating event camera technology with the era of Generative AI. We present an approach to address the key challenges in adapting video diffusion models to EVFI. We use data-efficient adaptation to bridge the gap in data quality and size between large-scale video generation datasets and smaller EVFI datasets, minimizing the risk of catastrophic forgetting. Our method also includes test-time optimization techniques: Per-tile Denoising and Fusion to reduce encoding/decoding losses between pixel and latent spaces, preserving high-fidelity appearance and accurate event-based motion control, and Two-side Fusion, which incorporates both start and end frame information at each denoising step for improved interpolation consistency. We also discuss current limitations in pre-trained video diffusion models for EVFI. Lastly, we demonstrate our model’s superior generalization and frame consistency on unseen real-world data, highlighting the motivation and potential of our approach.
7 Acknowledgment
We greatly acknowledge NSF’s support under awards OISE 2020624 and BCS 2318255
Supplementary Material
8 Video Results
Please refer to our project page: https://vdm-evfi.github.io/ for video results, which clearly demonstrate that our reconstructions provide superior consistency and generalization compared to other baselines.
9 Video Generation Task Results
As explained in the main paper, our method supports Event-based Video Generation, an extrapolation task that relies on only one frame (start or end) and events, unlike interpolation, which uses both frames. This constraint in video generation leads to error accumulation in the generated video, as shown in the last video of the website. We present a comparison of video generation and interpolation results on the BS-ERGB dataset, for video generation, only the left-end frame next to skip frames and corresponding events are used to generate the skipped frames, as shown in Table 3, relying on information from only one side instead of both start and end frames causes the PSNR and SSIM metrics to drop significantly compared to the interpolation results.
Task BS-ERGB (3 skips)
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Video Generation 25.79 0.84 0.11
Video Interpolation 27.74 0.88 0.12
Table 3: Comparison of our pipeline’s performance on Video Generation and Video Interpolation tasks on the BS-ERGB dataset.
We also show a qualitative comparison in Figure 9, as we can see because the video generation task is a extrapolation task that only uses the information from one side (start or end frame) of frames and events, it cannot avoid hallucination for the occluded/missing parts that are not present in the condition image. In contrast, the video generation task can mitigate the hallucination well by using complementary information from both frames.
Figure 9: An example illustrating the difference between event-based video generation, which relies solely on the start frame, and event-based video interpolation, which uses both start and end frames to infer the interpolated frames. We present the 3rd, 6th, and 9th interpolated frames for the 11 skips interpolation between start and end frames on Clear-Motion test sequences. In the video generation scenario, the model hallucinates the occluded parts behind the paper due to the lack of information in the start frame. In contrast, video interpolation avoids hallucination as the end frame provides the necessary information.
In addition to hallucination, relying on frame information from only one side leads to error accumulation during generation, as shown in Figure 10, the results in video generation (incorrect color accumulation on the finger) are inconsistent with the information provided by the frame on the other side. In contrast, interpolation ensures the reconstructed video remains consistent with the information from both frames. The comparison of video consistency is best observed in the last video of the website, which compares event-based video generation with event-based video interpolation.
Figure 10: An example illustrating the difference in video consistency between event-based video generation and interpolation for 11 skips between start and end frames on Clear-Motion test sequences, we present the 5th, 8th, and 11th interpolated frames. In the video generation task, color errors on the finger accumulate over time, with the 11th frame (the final interpolated frame) failing to align with the information in the end frame. In contrast, interpolation reduces error accumulation and ensures consistency by leveraging information from both the start and end frames.
10 Clear-Motion Test Sequences
To robustly evaluate the zero-shot generalization performance of all models on unseen real-world event-based video frame interpolation scenarios, we collected the Clear-Motion Test Sequences solely for testing purposes.
10.1 Event-RGB Aligned Video Capture Setup
In this section, we will present the capture setup for our event-rgb aligned video sequences, as shown in Figure 11.
Figure 11: An illustration of our capture setup for the Clear-Motion test sequences of Event-RGB aligned video sequences. The setup consists of three main components: an event camera, an RGB camera, and a beam splitter to align the field of view for both cameras.
We use the Prophesee EVK4 HD as our event camera, offering a capture resolution of 1280 $\times$ 720. For the RGB camera, we use the BFS-U3-31S4C-C Blackfly S, which provides a resolution of 2048 $\times$ 1536 and supports up to 55 frames per second (fps). To align the field of view for both cameras, we utilize the Thorlabs CCM1-BS013 30 mm Cage Cube-Mounted Non-Polarizing Beam Splitter. Additionally, we perform spatial and temporal alignment to synchronize the events with the captured RGB frames. After the spatial alighment, our final captured RGB frames aligned with event are of resolution 940 $\times$ 720 and 40 fps.
10.2 Details of Data Sequence
The Clear-Motion test sequences are designed to include clear and large motions, encompassing both camera and object movements, as well as objects and motion patterns distinct from those found in most existing real-world Event-based Video Frame Interpolation (EVFI) datasets [41, 21, 39]. This setup enables a straightforward evaluation of the generalization and consistency of various video frame interpolation methods. Table 4 provides details of our collected test sequences, including explanations for each sequence. Category (i) represents sequences with object motion, while Category (ii) represents sequences with camera motion. We also include a Figure 12 to show some example data in our test sequences.
Figure 12: An illustration of example frames overlaid with events from Clear-Motion test sequences.
Sequence #Frames Explanation Category
Paper_Shifting 200 The translational and 3D rotational motion of a paper with a simple texture (i)
Paper_Waving 200 The waving and 3D rotational motion of a paper with a simple texture (i)
Paper_Deforming 200 The deformation motion of a paper with a simple texture (i)
Camera_Far 200 The moving cameras capturing distant objects (ii)
Camera_Close 200 The moving cameras capturing nearby objects (ii)
Checkerboard_Planar 200 The planar translation and rotation of a nearby dense checkerboard (i)
Checkerboard_Depth 200 The motion of a checkerboard along the depth direction (i)
Checkerboard_3D 200 The 3D translation and rotation of a nearby dense checkerboard (i)
Texture_Box 200 The translation and rotation of a nearby highly textured box (i)
Table 4: A detailed description of our collected Clear-Motion test sequences, with sequence name, number of frames and explanation of each sequence. Category (i) includes sequences with object motion, while Category (ii) includes sequences with camera motion.
11 The Impact of Input Upsampling
As discussed in the main paper, to mitigate the loss of appearance and motion control accuracy caused by the conversion between downsampled latent space and pixel space in Latent Diffusion Models (LDM) [7], we employ test-time optimization in the Per-tile Denoising and Fusion process. This involves first upsampling the input image and event representations by a specified factor and then breaking them into fixed-size overlapping tiles before feeding them into the video diffusion process. The performance comparison across different upsampling factors is shown in Table 5. As the upsampling factor increases from 1 to 2, our model’s performance improves significantly, with PSNR increasing by approximately 3 dB, SSIM by 0.11, and LPIPS decreasing by 0.03. These results demonstrate the effectiveness of upsampling in the Per-tile Denoising and Fusion process, enhancing both the details in reconstructed frames and the accuracy of event-based motion control.
Method BS-ERGB (3 skips)
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Ours_1 24.82 0.77 0.15
Ours_1.5 25.86 0.82 0.16
Ours_2 27.74 0.88 0.12
[…]
To qualitatively assess the effect of upsampling, Figure 13 shows that as the upsampling factor increases from 1 to 2, details on the human eyes and fingers (e.g., nails and textures) improve significantly. Both quantitative and qualitative results highlight the effectiveness of upsampling in the Per-tile Denoising and Fusion process, enhancing the realism and event-based motion control accuracy of video interpolation results.
Figure 13: An illustration showcasing the qualitative impact of input upsampling factors. The leftmost image is the reference frame. (a) shows the interpolated result with an upsampling factor of 1, (b) with a factor of 1.5, and (c) with a factor of 2. As the upsampling factor increases from 1 to 2, the details on the human eyes and fingers, highlighted by red arrows, improve significantly.
12 More Visual Results
In this section, we provide additional visual results showcasing qualitative comparisons between our method and the baselines, as shown in Figures 14, 15, 16, 17, 18, and 19.
13 Additional Implementation Details
In this section, we provide additional implementation details. The pre-trained video diffusion model we used is Stable Video Diffusion [7] for 14-frame image-to-video generation. We trained our model with an effective batch size of 64, using a batch size of 4 per GPU and a gradient accumulation factor of 16. Training was conducted solely on the BS-ERGB dataset, and the model was tested on other unseen datasets without fine-tuning. All training was performed on 4 NVIDIA RTX A6000 GPUs, each with 50GB of memory. For training, we use the AdamW optimizer [24] with a learning rate of $5\times 10^{-5}$ and parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=1\times 10^{-8}$, and a weight decay of $1\times 10^{-2}$. Our model is trained on the BS-ERGB dataset for 72 hours to denoise noisy video latents for 3 and 11 skipped frames. All testing and inference on unseen data or datasets are conducted without fine-tuning, using the checkpoint trained on the BS-ERGB dataset.
During training and testing, the input to our adapted video diffusion model consists of 512 × 320 size tiles. Our model is trained solely to denoise/generate video latents using start frames and forward-time events. For testing and inference, we use an overlapping ratio of 0.1 for overlapping tiles and set the number of denoising steps in the video diffusion process to 25. The Per-tile Denoising and Fusion (for high-resolution frame reconstruction and event-based motion control) and Two-side Fusion (for converting video generation to interpolation) are both test-time optimization processes that do not require additional training.
14 Model Run Time, Memory, and Parameter Comparison
Table 6 reports testing results for all models run on a single NVIDIA RTX 4090 GPU. Each method generated $1024\times 576$ frames with run time averaged over 16 frames. VDM based methods (Time-Reversal, DynamiCrafter, and Ours) are more memory-intensive and time-consuming than other methods.
Method Run Time (s) Memory Usage (GB) #Parameters (M)
RIFE 0.5 0.9 9.8
CBMNet-Large 41.7 17.8 22.2
Time-Reversal 62.6 21.5 1524.6
PerVFI 9.3 5.2 13.9
InterpAny-Clearer 0.5 1.1 10.7
DynamiCrafter 92.0 18.3 1438.9
EMA-VFI 1.8 4.8 65.7
GIMM-VFI 3.0 9.5 19.8
Ours 200.1 17.2 2206.8
Table 6: Model Run Time, Memory, and Parameter comparison.
Figure 14: Additional baseline results on the Clear-Motion sequence Paper_Waving.
Figure 15: Additional baseline results on the BS-ERGB sequence as presented in the main paper.
Figure 16: An illustration showcasing the qualitative comparison on the Clear-Motion sequence Camera_Far, which involves large camera motion capturing distant objects, with 11 skips between the start and end frames. We present the interpolated 4th, 7th, and 10th frames. (Zoom in for the best viewing experience)
Figure 17: An illustration showcasing the qualitative comparison on the Clear-Motion sequence Camera_Close, which involves large camera motion capturing nearby objects, with 11 skips between the start and end frames. We present the interpolated 4th, 7th, and 10th frames. (Zoom in for the best viewing experience)
Figure 18: An illustration showcasing the qualitative comparison on the Clear-Motion sequence Checkerboard_Planar, which involves large planar motion of a nearby checkerboard, with 11 skips between the start and end frames. We present the interpolated 4th, 7th, and 10th frames. (Zoom in for the best viewing experience)
Figure 19: An illustration showcasing the qualitative comparison on the HQF dataset for the sequence poster_pillar_1, involving moving cameras capturing nearby posters, with 3 skips between the start and end frames. All interpolated frames are presented. (Zoom in for the best viewing experience)